# The Leftmost particle of branching subordinators

### v1

@@ Abstract
We define a family of continuous-time branching particle systems on the non-negative real line called branching subordinators and closely related to branching Lévy processes introduced by Bertoin and Mallein [13]. We pay a particular attention to the asymptotic behavior of the leftmost particle of branching subordinators and it turns out, under some assumptions, that the growth of the position of this leftmost particle is sub-linear thus making a significant change compared with the usual linear growth classically observed in the case of regular branching random walks with non-negative displacement distribution.
60J80,
branching subordinators,
branching Lévy processes,
subordinators,
leftmost position,
@@ 1 Introduction
We consider continuous-time particle systems on the non-negative real line called branching subordinators where particles move as independent subordinators. Each particle can also split (at possibly infinite rate) into several children (possibly infinitely many) whose positions relatively to the position of the parent are random. Branching subordinators can be seen as age-dependent branching processes (see [17] for example) and are in the continuity of branching Lévy processes introduced by Bertoin and Mallein [13]. One can think of branching Lévy processes as continuous-time counterparts of branching random walks. They generalize very famous processes such as branching Brownian motion, continuous-time branching random walks (see [23] for instance) and processes bearing the same name, that is branching Lévy processes in the sense of Kyprianou [20]. General branching Lévy processes first appeared in a work of Bertoin [9] in a less general setting than in [13] as a tool to study fragmentation processes, see the end of this introduction for details. Branching Lévy processes are also fundamental in the construction of Self-Similar Markov trees introduced by Bertoin, Curien and Riera in the book [3]. Roughly, the latter objects are related to branching Lévy processes in the same way positive self-similar Markov processes are related to classical Lévy processes via Lamperti representation.
@@
Introduce $\mathscr{Q}$, the space of point measures $\bm{\mu}$ on $[0,\infty)$ that are such that $\bm{\mu}([0,a])\in\mathbb{Z}_{+}$ for any $a\geq 0$. Each element of $\mathscr{Q}$ can be identified with a non-decreasing sequence $\bm{x}=(x_{n})_{n\geq 1}$ on $[0,\infty]$ such that $\lim_{n\to\infty}x_{n}=\infty$. The identification consists in setting $x_{n+1}=\ldots=x_{n+k}=a$ if $\bm{\mu}([0,a])=n$ and $\bm{\mu}(\{a\})=k$ (for $n\geq 0$, $k\geq 1$, $a\in[0,\infty)$), and for any $n\geq 1$, $x_{n}=\infty$ if $\bm{\mu}([0,\infty))<n$. We denote :=(,$\emptyset$:=($\infty$,$\ldots$)∅ := ( ∞ , … ) and $(0,\emptyset):=(0,\infty,\ldots)$. We equip $\mathscr{Q}$ with the vague topology: we say that a sequence $((y^{n}_{1},y^{n}_{2},\ldots))_{n}$ of elements of $\mathscr{Q}$ converges to $(y_{1},y_{2},\ldots)\in\mathscr{Q}$ if and only if, for any continuous function $f:[0,\infty]\to\mathbb{R}$ with compact support on $[0,\infty)$ and $f(\infty)=0$
@@
with $\mathscr{Q}^{\star}:=\{\bm{x}\in\mathscr{Q},x_{2}<\infty\}$. The measure $\Lambda$ encodes for both the structure of jumps and the structure of branchings in the associated particle system, while $d$ is the speed of the continuous component of the motion of particles. More precisely, when $\Lambda(\mathscr{Q}^{\star})<\infty$ (referred to as the case of branching subordinators with finite birth intensity in [13]), the dynamic of a branching subordinator with characteristics $(d,\Lambda)$ can be described as follows: there is initially one particle in the system denoted by $\varnothing$ and, during its lifetime, the particle $\varnothing$ moves and experiences branchings. Before the first branching occurs, $\varnothing$ moves according to a subordinator $\xi$ with drift $d$ and Lévy measure the image of $\Lambda(\cdot\cap(\mathscr{Q}\setminus\mathscr{Q}^{\star}))$ by the projection $\bm{x}=(x_{n})_{n\geq 1}\mapsto x_{1}$. Let $(T,\bm{X})=(T,(X_{n})_{n\geq 1})$ be a $(0,\infty)\times\mathscr{Q}^{\star}$-valued random variable with law $\Lambda(\mathscr{Q}^{\star})e^{-t\Lambda(\mathscr{Q}^{\star})}\mathrm{d}t\otimes\Lambda(\cdot\cap\mathscr{Q}^{\star})/\Lambda(\mathscr{Q}^{\star})$ and independent of $\xi$. At time $T$, the first branching occurs: the particle $\varnothing$ makes a jump of size $X_{1}$ and simultaneously gives progeny to a (possibly infinite) set of children such that at birth, the $i$-th offspring exists if and only if $X_{i+1}<\infty$ and is then located at $\xi(T-)+X_{i+1}$, $\xi(T-)$ denoting the pre-jump position of $\varnothing$. Starting from $\xi(T-)+X_{1}$, the particle $\varnothing$ performs according to an independent copy of $\xi$ until the next branching event occurs, according to an independent copy of $(T,\bm{X})$, and so on. In turn, each newborn particle moves and experiences branchings as described previously for $\varnothing$, independently of other particles in the system. Note that the assumption $\Lambda(\{(0,\emptyset)\})=0$ removes transitions that have no effect while $\Lambda(\{\emptyset\})=0$ says that there is no killing. The second part of (3) ensures the summability of small jumps and is analogous to the integrability condition satisfied by Lévy measures of classical subordinators. If $\Lambda(\mathscr{Q}^{\star})=0$, then there is no branching so we have a classical subordinator. Eliminating this degenerate case is the reason for assuming the third part of (3). Therefore, a particle never dies and we obtain a system of persisting particles such that the full trajectory of any particle is a subordinator with drift $d$ and Lévy measure $\Lambda_{1}$, the image of $\Lambda$ by the projection $\bm{x}=(x_{n})_{n\geq 1}\mapsto x_{1}$. These trajectories will be referred to as the canonical trajectories in the following, see Definition A.2.
@@
$\forall a>0,\int_{\mathscr{Q}}\big{|}\sharp\{n\geq 1:\;x_{n}<a\}-1\big{|}\Lambda(\mathrm{d}\bm{x})<\infty,$ (4)
@@
$\forall a>0,\int_{\mathscr{Q}}\big{|}\sharp\{n\geq 1;\;x_{n}<a\}-1\big{|}^{2}\Lambda(\mathrm{d}\bm{x})$ <<$\infty$.< ∞ . (5)
We construct the branching subordinator with characteristic couple $(d,\Lambda)$ for any $(d,\Lambda)$ with $d\geq 0$ and $\Lambda$ satisfying (3) and (4) in Appendix A. We see that thanks to equation (4), whether $\Lambda(\mathscr{Q}^{\star})$ is finite or not, $\mathscr{Q}^{\star}=\cup_{a>0}\{\bm{x}\in\mathscr{Q};\;x_{2}<a\}$ and $\Lambda(\{\bm{x}\in\mathscr{Q};\;x_{2}<a\})<\infty$ for any $a>0$ and this will be crucial in the construction. Let us stress on the fact that arguments presented in Appendix A are borrowed from the construction of branching Lévy processes of Bertoin and Mallein [13] and Shi and Watson [22], but a slight adaptation is required in our case since we make weaker hypothesis. To index the set of particles in a branching subordinator $Y$, we use a generalization of the Ulam–Harris notation, introduced by Shi and Watson [22]. For any $t\geq 0$, we then denote by $\mathscr{N}(t)$ the set of particles in $Y$ alive at time $t$ and for any particle $u\in\mathscr{N}(t)$ and $s\leq t$, $Y_{u}(s)$ denotes the position of $u$ at time $s$, if already born at time $s$. Otherwise, $Y_{u}(s)$ stands for the position of the most recent ancestor of $u$ alive at time $s$, see Appendix A for a rigorous definition.
In general, (4) is not sufficient to ensure the well-definiteness of a branching Lévy process. The fact that in our case, trajectories are non-decreasing is crucial. Instead of (4), it is usually required (see equation (1.4) in [13]) that there exists $\theta\geq 0$ such that
@@
where, by convention, $e^{-\theta\times\infty}=0$ for any $\theta\geq 0$. If $\mathscr{E}(\Lambda):=\{\lambda\geq 0;\;\int_{\mathscr{Q}}\left(\sum_{k=2}^{\infty}e^{-\lambda x_{k}}\right)\Lambda(\mathrm{d}\bm{x})<\infty\}$, then (6) implies that $[\theta,\infty)\subset\mathscr{E}(\Lambda)$. For any $\lambda\in\mathscr{E}(\Lambda)$, we define $\kappa(\lambda)$ by
@@
and $\kappa(\lambda)=-\infty$ otherwise. When (6) holds, then, by a slight adaptation of Theorem 1.1(ii) in [13], a branching subordinator $Y=(Y_{u}(t);u\in\mathscr{N}(t),\;t\geq 0)$ with characteristics $(d,\Lambda)$ satisfies, for all $t\geq 0$ and any $\lambda\in\mathscr{E}(\Lambda)$
@@
In that case, if, for any $\lambda\in\mathscr{E}(\Lambda)$, we denote $W_{\lambda}(t):=e^{t\kappa(\lambda)}\sum_{u\in\mathscr{N}(t)}e^{-\lambda Y_{u}(t)}$, then $(W_{\lambda}(t))_{t\geq 0}$ is a non-negative martingale thus converging to a random variable $W_{\infty}(\lambda)\in[0,\infty)$ as $t\to\infty$. Bertoin and Mallein proved (Theorem 1.1 in [12]) that $(W_{\lambda}(t))_{t\geq 0}$ is uniformly integrable if and only if $\lambda\kappa^{\prime}(\lambda)>\kappa(\lambda)$ and $\int_{\mathscr{Q}}\sum_{k\geq 1}e^{-\lambda x_{k}}(\log\sum_{k\geq 1}e^{-\lambda x_{k}}-1)^{+}\Lambda(\mathrm{d}\bm{x})<\infty$, with $\mathfrak{z}^{+}=\max(\mathfrak{z},0)$. Otherwise, $W_{\infty}(\lambda)=0$ almost surely. This is a slightly different version of the Biggin’s martingale convergence theorem (originally dedicated to branching random walks, see for instance [11] and [2]) in branching Lévy processes settings.
@@
As mentioned at the very beginning of the Introduction, branching Lévy processes are closely related to fragmentation processes. The latter processes were initially studied by Kolmogorov [19] and describe the evolution of particles randomly breaking over time. Each particle is characterized by its mass. Let us mention in particular homogeneous fragmentation [5] (see also [8, Sec. 3.1]) and its generalization, the compensated fragmentation introduced by Bertoin [9]. The homogeneous fragmentation is related to branching subordinators while the compensated fragmentation is related to branching spectrally negative Lévy processes. Informally, homogeneous fragmentation can be described as follows: a particle of initial mass $1$ is possibly subject to continuous erosion and breaks (at possibly infinite rate) into children particles (possibly infinitely many). At a splitting time, a particle gives progeny to a set of children with birth mass $\mathfrak{m}p_{1},\mathfrak{m}p_{2},\cdots$ where $\sum_{i\geq 1}p_{i}\leq 1$ and $\mathfrak{m}$ is the mass of the particle just before splitting. Each newborn particle behaves similarly as its parent and independently of other particle in the system and so on. Let $Y$ be a branching subordinator with characteristics $(d,\Lambda)$ where $\Lambda$ satisfies (3) and $\Lambda(\{\bm{x}\in\mathscr{Q},\sum_{i\geq 1}e^{-x_{i}}>1\})=0$ (this implies (6) since we then have $\int_{\mathscr{Q}}(\sum_{k=2}^{\infty}e^{-\theta x_{k}})\Lambda(\mathrm{d}\bm{x})\leq\int_{\mathscr{Q}}(1-e^{-x_{1}})\Lambda(\mathrm{d}\bm{x})<\infty$). Then one can see that the process $(e^{-Y_{u}(t)};\;u\in\mathscr{N}(t),\;t\geq 0)$ is a homogeneous fragmentation process with erosion coefficient $d$ and whose dislocation measure is the image measure of $\Lambda$ by $\bm{x}\mapsto(e^{-x_{1}},e^{-x_{2}},\ldots)$. Here, $e^{-Y_{u}(t)}$ denotes the mass of a particle $u$ alive at time $t$. We also refer to [14] and [22] for more details on links between fragmentation processes and branching Lévy processes. Note that there exists another type of fragmentation called self-similar fragmentation for which homogeneous fragmentation corresponds to the special case of fragmentation with index of self-similarity 00, see [6].
@@
One can notice that the largest particle of a homogeneous fragmentation process and the position of the leftmost particle of the associated branching subordinator are directly related: $\max_{u\in\mathscr{N}(t)}e^{-Y_{u}(t)}=e^{-\underline{Y}(t)}$. The behavior of the largest particle, or largest fragment, has been deeply studied in the past few years, see for instance [7] for the homogeneous fragmentation and [14], [15] for self-similar fragmentation.

### v2

@@ Abstract
We define a family of continuous-time branching particle systems on the non-negative real line, called branching subordinators, where particles move as independent subordinators. Each particle can also split (at possibly infinite rate) into several children (possibly infinitely many) whose positions relative to the position of the parent are random. These particle systems are in the continuity of branching Lévy processes introduced by Bertoin and Mallein [14]. We pay a particular attention to the asymptotic behavior of the leftmost particle of branching subordinators. It turns out that, under some assumptions, the rate of growth of the position of the leftmost particle is of order $t^{\gamma}$ where $\gamma\in(0,1)$ depends explicitly on the parameters. This sub-linear growth is significantly different from the classical linear growth observed for regular branching random walks with non-negative displacements.
@@ 1 Introduction
We consider continuous-time particle systems on the non-negative real line called branching subordinators where particles move as independent subordinators. Each particle can also split (at possibly infinite rate) into several children (possibly infinitely many) whose positions relatively to the position of the parent are random. Branching subordinators can be seen as age-dependent branching processes, see for example [19], and the end of this introduction for details. They are in the continuity of branching Lévy processes introduced by Bertoin and Mallein [14]. One can think of branching Lévy processes as continuous-time counterparts of branching random walks. They generalize very famous processes such as branching Brownian motion, continuous-time branching random walks (see [27] for instance) and processes bearing the same name, that is branching Lévy processes in the sense of Kyprianou [22]. General branching Lévy processes first appeared in a work of Bertoin [10] in a less general setting than in [14] as a tool to study fragmentation processes, see the end of this introduction for details. Branching Lévy processes are also fundamental in the construction of Self-Similar Markov trees introduced by Bertoin, Curien and Riera in the book [4]. Roughly, the latter objects are related to branching Lévy processes in the same way positive self-similar Markov processes are related to classical Lévy processes via Lamperti representation.
@@
Introduce $\mathscr{Q}$, the space of point measures $\bm{\mu}$ on $0,\infty)$ that are such that $\bm{\mu}([0,a])\in\mathbb{Z}_{+}$ for any $a\geq 0$. Each element of $\mathscr{Q}$ can be identified with a non-decreasing sequence $\bm{x}=(x_{n})_{n\geq 1}$ on $0,\infty]$ such that $\lim_{n\to\infty}x_{n}=\infty$. The identification consists in setting $x_{n+1}=\ldots=x_{n+k}=a$ if $\bm{\mu}([0,a))=n$ and $\bm{\mu}(\{a\})=k$ (for $n\geq 0$, $k\geq 1$, $a\in[0,\infty)$), and for any $n\geq 1$, $x_{n}=\infty$ if $\bm{\mu}([0,\infty))<n$. We denote $\emptyset:=(\infty,\ldots)$ and $(0,\emptyset):=(0,\infty,\ldots)$. We equip $\mathscr{Q}$ with the vague topology: we say that a sequence $((y^{n}_{1},y^{n}_{2},\ldots))_{n}$ of elements of $\mathscr{Q}$ converges to $(y_{1},y_{2},\ldots)\in\mathscr{Q}$ if and only if, for any continuous function $f:[0,\infty]\to\mathbb{R}$ with compact support on $0,\infty)$ and $f(\infty)=0$
@@
with $\mathscr{Q}^{\star}:=\{\bm{x}\in\mathscr{Q};x_{2}<\infty\}$. The measure $\Lambda$ encodes for both the structure of jumps and the structure of branchings in the associated particle system, while $d$ is the speed of the continuous component of the motion of particles. More precisely, when $\Lambda(\mathscr{Q}^{\star})<\infty$ (referred to as the case of branching subordinators with finite birth intensity in [14]), the dynamic of a branching subordinator with characteristics $(d,\Lambda)$ can be described as follows: there is initially one particle in the system denoted by $\varnothing$ and, during its lifetime, the particle $\varnothing$ moves and experiences branchings. Before the first branching occurs, $\varnothing$ moves according to a subordinator $\xi$ with drift $d$ and Lévy measure the image of $\Lambda(\cdot\cap(\mathscr{Q}\setminus\mathscr{Q}^{\star}))$ by the projection $\bm{x}=(x_{n})_{n\geq 1}\mapsto x_{1}$. Let $(T,\bm{X})=(T,(X_{n})_{n\geq 1})$ be a $(0,\infty)\times\mathscr{Q}^{\star}$-valued random variable with law $\Lambda(\mathscr{Q}^{\star})e^{-t\Lambda(\mathscr{Q}^{\star})}\mathrm{d}t\otimes\Lambda(\cdot\cap\mathscr{Q}^{\star})/\Lambda(\mathscr{Q}^{\star})$ and independent of $\xi$. At time $T$, the first branching occurs: the particle $\varnothing$ makes a jump of size $X_{1}$ and simultaneously gives progeny to a (possibly infinite) set of children such that, at birth, the $i$-th offspring exists if and only if $X_{i+1}<\infty$ and is then located at $\xi(T-)+X_{i+1}$, $\xi(T-)$ denoting the pre-jump position of $\varnothing$. Starting from $\xi(T-)+X_{1}$, the particle $\varnothing$ performs according to an independent copy of $\xi$ until the next branching event occurs, according to an independent copy of $(T,\bm{X})$, and so on. In turn, each newborn particle moves and experiences branchings as described previously for $\varnothing$, independently of other particles in the system. Note that the assumption $\Lambda(\{(0,\emptyset)\})=0$ removes transitions that have no effect while $\Lambda(\{\emptyset\})=0$ says that there is no killing. The second part of (3) ensures the summability of small jumps and is analogue to the integrability condition satisfied by Lévy measures of classical subordinators. If $\Lambda(\mathscr{Q}^{\star})=0$, then there is no branching so we have a classical subordinator. Eliminating this degenerate case is the reason for assuming the third part of (3). Therefore, a particle never dies and we obtain a system of persisting particles such that the full trajectory of any particle is a subordinator with drift $d$ and Lévy measure $\Lambda_{1}$, the image of $\Lambda$ by the projection $\bm{x}=(x_{n})_{n\geq 1}\mapsto x_{1}$. These trajectories will be referred to as the canonical trajectories in the following, see Definition A.2.
@@
$\forall a>0,\int_{\mathscr{Q}}\big|\sharp\{n\geq 1;\;x_{n}<a\}-1\big|\Lambda(\mathrm{d}\bm{x})<\infty,$ (4)
@@
$\forall a>0,\int_{\mathscr{Q}}\big|\sharp\{n\geq 1;\;x_{n}<a\}-1\big|^{2}\Lambda(\mathrm{d}\bm{x})$ $<\infty.$ (5)
We construct the branching subordinator with characteristic couple $(d,\Lambda)$ for any $(d,\Lambda)$ with $d\geq 0$ and $\Lambda$ satisfying (3) and (4) in Appendix A. We see that thanks to equation (4), whether $\Lambda(\mathscr{Q}^{\star})$ is finite or not, $\mathscr{Q}^{\star}=\cup_{a>0}\{\bm{x}\in\mathscr{Q};\;x_{2}<a\}$ and $\Lambda(\{\bm{x}\in\mathscr{Q};\;x_{2}<a\})<\infty$ for any $a>0$ and this will be crucial in the construction. Let us stress on the fact that arguments presented in Appendix A are borrowed from the construction of branching Lévy processes of Bertoin and Mallein [14] and Shi and Watson [26], but we provide a slight adaptation tailored to our case. To index the set of particles in a branching subordinator $Y$, we use a generalization of the Ulam–Harris notation, introduced by Shi and Watson [26]. For any $t\geq 0$, we then denote by $\mathscr{N}(t)$ the set of particles in $Y$ alive at time $t$ and for any particle $u\in\mathscr{N}(t)$ and $s\leq t$, $Y_{u}(s)$ denotes the position of $u$ at time $s$, if already born at time $s$. Otherwise, $Y_{u}(s)$ stands for the position of the most recent ancestor of $u$ alive at time $s$, see Appendix A for a rigorous definition.
In general, (4) is not sufficient to ensure the well-definiteness of a branching Lévy process. The fact that, in our case, trajectories are non-decreasing is crucial. Instead of (4), it is usually required (see equation (1.4) in [14]) that there exists $\theta\geq 0$ such that
@@
where, by convention, $e^{-\theta\times\infty}=0$ for any $\theta\geq 0$. If $\mathscr{E}(\Lambda):=\{\lambda\geq 0;\;\int_{\mathscr{Q}}\left(\sum_{k=2}^{\infty}e^{-\lambda x_{k}}\right)\Lambda(\mathrm{d}\bm{x})<\infty\}$, then (6) implies that $\theta,\infty)\subset\mathscr{E}(\Lambda)$. For any $\lambda\in\mathscr{E}(\Lambda)$, we define $\kappa(\lambda)$ by
@@
and $\kappa(\lambda)=$-$\infty$ otherwise. When (6) holds, then, by a slight adaptation of Theorem 1.1(ii) in [14], a branching subordinator $Y=(Y_{u}(t);u\in\mathscr{N}(t),\;t\geq 0)$ with characteristics $(d,\Lambda)$ satisfies, for all $t\geq 0$ and any $\lambda\in\mathscr{E}(\Lambda)$,
@@
In that case, if, for any $\lambda\in\mathscr{E}(\Lambda)$, we denote
$W_{\lambda}(t):=e^{t\kappa(\lambda)}\sum_{u\in\mathscr{N}(t)}e^{-\lambda Y_{u}(t)},$ (8)
then $(W_{\lambda}(t))_{t\geq 0}$ is a non-negative martingale thus converging to a random variable $W_{\lambda}(\infty)\in[0,\infty)$ as $t\to\infty$. Bertoin and Mallein proved (Theorem 1.1 in [13]) that $(W_{\lambda}(t))_{t\geq 0}$ is uniformly integrable if and only if $\lambda\kappa^{\prime}(\lambda)>\kappa(\lambda)$ and $\int_{\mathscr{Q}}\sum_{k\geq 1}e^{-\lambda x_{k}}(\log\sum_{k\geq 1}e^{-\lambda x_{k}}-1)^{+}\Lambda(\mathrm{d}\bm{x})<\infty$, with $\mathfrak{z}^{+}=\max(\mathfrak{z},0)$. Otherwise, $W_{\lambda}(\infty)=0$ almost surely. This is a slightly different version of the Biggin’s martingale convergence theorem (originally dedicated to branching random walks, see for instance [12] and [2]) in branching Lévy processes settings. Another important martingale for branching random walks known as the derivative martingale has been studied in the context of branching Lévy processes, see [24].
@@
As mentioned at the very beginning of the Introduction, branching Lévy processes are closely related to fragmentation processes. The latter processes were initially studied by Kolmogorov [21] and describe the evolution of particles randomly breaking over time. Each particle is characterized by its mass. Let us mention in particular homogeneous fragmentation [6] (see also [9, Sec. 3.1]) and its generalization, the compensated fragmentation introduced by Bertoin [10]. The homogeneous fragmentation is related to branching subordinators while the compensated fragmentation is related to branching spectrally negative Lévy processes. Informally, homogeneous fragmentation can be described as follows: a particle of initial mass $1$ is possibly subject to continuous erosion and breaks (at possibly infinite rate) into children particles (possibly infinitely many). At a splitting time, a particle gives progeny to a set of children with birth mass $\mathfrak{m}p_{1},\mathfrak{m}p_{2},\cdots$ where $\sum_{i\geq 1}p_{i}\leq 1$ and $\mathfrak{m}$ is the mass of the particle just before splitting. Each newborn particle behaves similarly as its parent and independently of other particle in the system and so on. Let $Y$ be a branching subordinator with characteristics $(d,\Lambda)$ where $\Lambda$ satisfies (3) and $\Lambda(\{\bm{x}\in\mathscr{Q};\sum_{i\geq 1}e^{-x_{i}}>1\})=0$ (this implies (6) since we then have $\int_{\mathscr{Q}}(\sum_{k=2}^{\infty}e^{-\theta x_{k}})\Lambda(\mathrm{d}\bm{x})\leq\int_{\mathscr{Q}}(1-e^{-x_{1}})\Lambda(\mathrm{d}\bm{x})<\infty$). Then one can see that the process $(e^{-Y_{u}(t)};\;u\in\mathscr{N}(t),\;t\geq 0)$ is a homogeneous fragmentation process with erosion coefficient $d$ and whose dislocation measure is the image measure of $\Lambda$ by $\bm{x}\mapsto(e^{-x_{1}},e^{-x_{2}},\ldots)$. Here, $e^{-Y_{u}(t)}$ denotes the mass of a particle $u$ alive at time $t$. We also refer to [15] and [26] for more details on links between fragmentation processes and branching Lévy processes. Note that there exists another type of fragmentation called self-similar fragmentation for which homogeneous fragmentation corresponds to the special case of fragmentation with index of self-similarity 0, see [7].
@@
One can notice that the largest particle of a homogeneous fragmentation process and the position of the leftmost particle of the associated branching subordinator are directly related: $\max_{u\in\mathscr{N}(t)}e^{-Y_{u}(t)}=e^{-\underline{Y}(t)}$. The behavior of the largest particle, or largest fragment, has been deeply studied in the past few years, see for instance [8] for the homogeneous fragmentation and [15], [16], [17] for self-similar fragmentation.
Another motivation to study branching subordinators comes from age-dependent branching processes, and our model can be seen as an extension of these branching processes, where the heigh of a particle at birth is interpreted as a birth time, the displacement of the newborn particle with respect to the parent being a delay between the reproduction time of the parent and the birth of the child, and the jumps of a particle between reproductions are interpreted as dormancy periods. In this interpretation, the original time of the process is an abstract local time allowing to run some underlying Poisson processes. We are particularly interested in determining possible behaviors in the case of non-explosion, as the explosion for heavy-tailed age-dependent branching processes has received attention by the past (see [1] and more recently in [3]). Here the term explosion is a terminology used in the study of age-dependent branching processes and means that infinitely many particles are born in finite time. The analogue, in our context, of that phenomenon is discussed in details in Section 1.3.1.